OpenStudy (anonymous):

Take any number (except for 1). Square that number and then subtract one. Divide by one less than your original number. Now subtract your original number. You reached 1 for an answer, didn’t you? How does this number game work? (Hint: Redo the number game using a variable instead of an actual number and rewrite the problem as one rational expression). How did the number game use the skill of simplifying rational expressions? Create your own number game using the rules of algebra.

OpenStudy (anonymous):

now i'm doing this sum in the order they're telling me, check it, let my number be x,so x^2-1=(x-1)(x+1) (x-1)(x+1)/((x-1)=(x+1) (x+1)-1=x. TADA! the same number !